OpenStudy (anonymous):

In how many way can a boat crew of eight women be arranged if three of the women can only row on the bow side and two others can only row on the stroke side?

OpenStudy (anonymous):

You have three options for the fourth bow side person, after that you know who is on what side. After that there are 4*3*2*1 possibilities to seat the women per side. So the final answer is 3*4*3*2*1*4*3*2*1=3*(4!)^2

OpenStudy (anonymous):

i dont understand where the 4!^2 came from

OpenStudy (anonymous):

Do you understand how I got to 3*4*3*2*1*4*3*2*1?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

what i got was 3!x(3!x2!x3!) but i was wrong lol

OpenStudy (anonymous):

Do you see that there are three candidates for the fourth person on the bow side?

OpenStudy (anonymous):

but only three can row on the bow side

OpenStudy (anonymous):

Did I misunderstand the question? there are four women on both sides right?

OpenStudy (anonymous):

no only three on the bow side and 2 on the stroke side

OpenStudy (anonymous):

But you have eight women?

OpenStudy (anonymous):

ur answers is right i think I'm very confessed

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

The way I read it is that there are three women that can only row on the bow side, two that can only row on the stroke side and three women can row on both sides.

OpenStudy (anonymous):

i see i misunderstood the question

OpenStudy (anonymous):

i still don't get it lol

OpenStudy (anonymous):

Do you understand the question?

OpenStudy (anonymous):

i thought it meant that three women can only row and two women can only stroke and the other three don't do anything so you have to find how many combinations can occur with eight women on the boat

OpenStudy (anonymous):

I understood it differently, but we can do it like this too.

OpenStudy (anonymous):

this is how i thought it should be done; RRRSSOOO where R- row, S-stroke and O is other

OpenStudy (anonymous):

so theres three combinations 3!

OpenStudy (anonymous):

then you times it by the possibilities it then becomes: 3!x(3!x2!x3!)

OpenStudy (anonymous):

but i think it should be : 4! x (3!x2!x3!) cause that gets me the answer at the back of the book

OpenStudy (anonymous):

I'm just not really sure why or if my method is even correct lol

OpenStudy (anonymous):

Let me think about this for a while.

OpenStudy (anonymous):

ok:)

OpenStudy (anonymous):

4! x(3!x2!x3!)=3*(4!)^2 So my first answer gives you the same answer as your book. So I'm thinking my first explanation of the question is the right one.

OpenStudy (anonymous):

there are four possibilities to seat the women on both sides thats why its 4!